# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ B.2 Proof of Proposition 2
Suppose $\widehat{C}_{\alpha}$ satisfies the strong selection-conditional coverage (3) for a class of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, it holds that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$. By definition, and since $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint, we have
@@
FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}(\widehat{{\mathcal{S}}}),\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}$ (26)
@@
$=\sum_{r=1}^{m}\sum_{\ell\in{\mathcal{L}}}\mathbb{E}\bigg{[}\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\cdot\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}.$ (27)
@@
$\sum_{\ell\in{\mathcal{L}}}\sum^{m}_{j=1}\frac{1}{r(\ell)}\mathbb{E}\big{[}\mathds{1}\big{\{}\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell},j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j};j,\mathfrak{S}_{\ell},\widehat{{\mathcal{S}}})\big{\}}\bigg{]}$ (28)
@@
(28)$=\sum_{\ell\in{\mathcal{L}}}\frac{\alpha}{r(\ell)}\mathbb{E}\Big{[}|\widehat{{\mathcal{S}}}|\cdot\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\Big{]}=\alpha\cdot\mathbb{P}\big{(}\widehat{{\mathcal{S}}}\neq\varnothing\big{)}\leq\alpha.$
B.3 Proof of Proposition 3
Consider the following example with $m=2$ test units and ${\mathcal{Y}}=\mathbb{R}$. Suppose $\alpha\in(0,1)$. The selection rule returns
@@ B.3 Proof of Proposition 3
Consider a prediction set $\widehat{C}_{\alpha}$ such that
$\widehat{C}_{\alpha}\big{(}X_{n+1};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{1,2\}\big{)}=\mathbb{R},\quad\widehat{C}_{\alpha}\big{(}X_{n+1};\{1\}\big{)}=\varnothing,\quad\widehat{C}_{\alpha}\big{(}X_{n+2};\{2\}\big{)}=\varnothing.$
Since the construction is symmetric in $1$ and $2$, we focus on $j=1$ without loss of generality.
@@
For the FCR, we have
@@
Therefore, the constructed prediction set $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage but violates the FCR control.
B.4 Proof of Theorem 1
Proof of (a).
To prove (11), it suffices to show that
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{j}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S}\Big{)}\geq 1-\alpha.$ (32)
Recall that ${\mathcal{D}}_{j}={\mathcal{D}}_{{\textnormal{calib}}}\cup\{Z_{n+j}\}$, and ${\mathcal{D}}_{j}^{c}=\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}}$ is the set of test data without the $j$-th one; ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$ and ${\mathcal{I}}_{j}^{c}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$. Let $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ be the unordered set of ${\mathcal{D}}_{j}$, and we denote the realized value of the unordered set as $[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$, i.e., $z_{i}$’s are fixed realized values of $Z_{i}$’s, but we do not know which values in $[d_{j}]$ correspond to the calibration set and which to the test point. We also write $z_{i}=(x_{i},y_{i})$ and $v_{i}=V(x_{i},y_{i})$.
The exchangeability of $\{Z_{i}\}_{i\in{\mathcal{I}}_{j}}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$ implies that, given knowledge of the unordered realized data values in $[d_{j}]$, the probability of any assignment of these values to data points is equal. That is, for any permutation $\pi$ of ${\mathcal{I}}_{j}$,
@@ B.4 Proof of Theorem 1
Next, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we define the following subset of ${\mathcal{I}}_{j}$:
@@
where $[D_{\textnormal{calib}}^{i}]=[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ is an unordered set, and $D_{\textnormal{test}}^{i}=(X_{n+1},\dots,x_{i},\dots,X_{n+m})$. Note that the collection $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, so $\widetilde{R}$ is also fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$.
Recall that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})=\{i\in{\mathcal{I}}_{\textnormal{calib}}:j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\}$. We let ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})$ if $j\notin\widehat{\mathcal{S}}$ and ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}$ if $j\in\widehat{\mathcal{S}}$. We claim that given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
@@
where repetition of elements is allowed on both sides. To see this, let $\pi$ denote the permutation on ${\mathcal{I}}_{j}$ such that $V_{i}=v_{\pi(i)}$ and $V_{n+j}=v_{\pi(n+j)}$. Then
@@
for all $i\in\{1,\dots,n\}$. Denoting $\widehat{\mathcal{S}}=\widehat{\mathcal{S}}^{{\textnormal{swap}(n+j,j)}}$ for notational simplicity, we have,
$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}$ =$\overset{\rm(a)}{=}\big{\{}V_{i}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}\in\mathfrak{S}\big{\}}$
=$\overset{\rm(b)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}},\,\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}\big{\}}$
@@
=$=\big{\{}v_{\ell}\colon\ell\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\ell}],D_{{\textnormal{test}}}^{\ell}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\pi(i)}],D_{\textnormal{test}}^{\pi(i)}\big{)}\in\mathfrak{S}\big{\}}$
@@
Above, step (a) follows from the definition of ^$\widehat{\mathcal{R}}_{n+j}^{+}$; steps (b) and (c) are due to (34), and step (d) follows from the definition of ${\widetilde{R}}$. We thus completed the proof of Eqn. (33).
Returning to the proof of (32), we have by Bayes’ rule that
@@
Again, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, the only randomness is the assignment of $[z_{1},\dots,z_{n},z_{n+j}]$ to $(Z_{1},\dots,Z_{n},Z_{n+j})$. By definition and the exchangeability of ${\mathcal{D}}_{j}$ given $\{X_{\ell}\}_{\ell\in{\mathcal{I}}_{j}^{c}}$, it is straightforward that
On the other hand, by (33),
@@
Since ${\widetilde{R}}$ is determistic conditional on $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
@@
Combining the above and (37), we have thus verified (35). Then by the tower property, we conclude (32), and thus (11).
When the elements in ${\widetilde{R}}$ are distinct, we additionally have the following determistic inequality following from the definition of the quantile function:
@@
Proof of (b).
To conveniently denote repeated values in $[z_{1},\dots,z_{n},z_{n+j}]$ and $[v_{1},\dots,v_{n},v_{n+j}]$ we let $[1],[2],\dots,[n],[n+j]$ be a permutation of ${\mathcal{I}}_{j}$ so that $v_{[1]}\leq v_{[2]}\leq\cdots\leq v_{[n]}\leq v_{[n+j]}$. Then $[d_{j}]$ can be equivalently represented by $\{z_{[i]}\}_{i\in{\mathcal{I}}_{j}}$, and for any $i\in{\mathcal{I}}_{j}$,
@@
Following the proof of part (a) up to Equation (33), we know that
@@
On the other hand, due to (33) and the fact that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}=\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})$. when $j\in\widehat{\mathcal{S}}$, for any $t\in(0,1)$,
$\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}<v_{i}\}+U_{j}\sum_{i\in\widetilde{R}}\mathds{1}\{V_{n+j}=v_{i}\}}{|\widetilde{R}|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
@@
where we used the take-out property of conditional probability/expectation in the first step; the probability in the second line is only with respect to $U_{j}\sim\textnormal{Unif}[0,1]$, and the last equality is a standard result that is straightforward to check with exchangeability (see, e.g., Vovk et al., (2005, Proposition 2.4)). Putting them togther and letting $t=1-\alpha$, we have

### v3

@@ D.2 Proof of Proposition 2
Suppose $\widehat{C}_{\alpha,n+j}^{(\ell)}$ satisfies the strong selection-conditional coverage (3) for a class of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, it holds that $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$. By definition, and since $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint, we have
@@
FCR $=\mathbb{E}\bigg{[}\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}$ (30)
@@
$=\sum_{\ell\in{\mathcal{L}}}\mathbb{E}\bigg{[}\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\cdot\frac{\sum_{j\in[m]}\mathds{1}\{j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}^{(\ell)}\}}{|\widehat{{\mathcal{S}}}|\vee 1}\bigg{]}.$ (31)
@@
$\sum_{\ell\in{\mathcal{L}}}\sum^{m}_{j=1}\frac{1}{r(\ell)\vee 1}\mathbb{E}\big{[}\mathds{1}\big{\{}\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell},j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha,n+j}^{(\ell)}\big{\}}\big{]}$
@@
(D.2)$=\sum_{\ell\in{\mathcal{L}}}\frac{\alpha}{r(\ell)\vee 1}\mathbb{E}\Big{[}|\widehat{{\mathcal{S}}}|\cdot\mathds{1}\{\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\}\Big{]}=\alpha\cdot\mathbb{P}\big{(}\widehat{{\mathcal{S}}}\neq\varnothing\big{)}\leq\alpha.$
D.3 Proof of Proposition 3
Consider the following example with $m=2$ test units and ${\mathcal{Y}}=\mathbb{R}$. Suppose the target miscoverage level $\alpha\in(0,1)$, and the selection rule returns
@@ D.3 Proof of Proposition 3
Consider prediction sets $\widehat{C}_{\alpha,n+1}$ and $\widehat{C}_{\alpha,n+2}$ such that
  • when $\widehat{{\mathcal{S}}}=\{1,2\}$, $\widehat{C}_{\alpha,n+1}=\mathbb{R},\widehat{C}_{\alpha,n+2}=\mathbb{R}$;
  • when $\widehat{{\mathcal{S}}}=\{1\}$, $\widehat{C}_{\alpha,n+1}=\varnothing$;
  • when $\widehat{{\mathcal{S}}}=\{2\}$, $\widehat{C}_{\alpha,n+2}=\varnothing$.
We can check that the weak selection-conditional coverage is satisfied for $j=1$:
@@
By symmetry, the weak selection-conditional coverage is also satisfied for $j=2$. Meanwhile, one can check that
@@
Therefore, the constructed prediction sets $\widehat{C}_{\alpha,n+1}$ and $\widehat{C}_{\alpha,n+2}$ satisfy the weak selection-conditional coverage but violate the FCR control.
D.4 Proof of Theorem 1
Recall that ${\mathcal{D}}_{j}={\mathcal{D}}_{{\textnormal{calib}}}\cup\{Z_{n+j}\}$ and ${\mathcal{D}}_{j}^{c}=\{X_{n+\ell}\}_{\ell\in[m]\backslash\{j\}}$. For notational convenience, we also denote the index sets ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$ and ${\mathcal{I}}_{j}^{c}=[m]\backslash\{j\}$. Let $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ be the unordered set of ${\mathcal{D}}_{j}$, and we denote the realized value of the unordered set as $[d_{j}]=[z_{1},\dots,z_{n},z_{n+j}]$, i.e., $z_{i}$’s are fixed realized values of $Z_{i}$’s, but we do not know which values in $[d_{j}]$ correspond to the calibration set and which to the test point. We also write $z_{i}=(x_{i},y_{i})$ and $v_{i}=V(x_{i},y_{i})$, with $x_{i}$ and $y_{i}$ being the realized values of $X_{i}$ and $Y_{i}$. We then prove the two statements separately.
Proof of (a). To prove (11), it suffices to show that
$\mathbb{P}\Big{(}V_{n+j}\leq{\textnormal{Quantile}}\big{(}1-\alpha;\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\cup\{+\infty\}\big{)}\,\Big{|}\,j\in\widehat{\mathcal{S}},\,\widehat{{\mathcal{S}}}\in\mathfrak{S}\Big{)}\geq 1-\alpha.$ (33)
The exchangeability of data in ${\mathcal{D}}_{j}$ given ${\mathcal{D}}_{j}^{c}$ implies that, given knowledge of the unordered realized data values in $[d_{j}]$, the probability of any assignment of these values to the data points is equal. That is, for any permutation $\pi$ of ${\mathcal{I}}_{j}$,
@@ D.4 Proof of Theorem 1
Next, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we define the following subset of ${\mathcal{I}}_{j}=[n]\cup\{n+j\}$:
@@
where $[D_{\textnormal{calib}}^{i}]=[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ is an unordered set ($[d_{j}]$ with $z_{i}$ removed) and $D_{\textnormal{test}}^{i}=(X_{n+1},\dots,x_{i},\dots,X_{n+m})$, i.e., the test data with $X_{n+j}$ replaced by $x_{i}$. Also, $[D_{\textnormal{calib}}^{n+j}]=[z_{1},\dots,z_{n}]$.
We note that a set is always unordered. By definition, the unordered collection of $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$. That is, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, no matter which value in $[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ the test point $Z_{n+j}$ takes on, the unordered set of $\{{\mathcal{S}}([D_{\textnormal{calib}}^{i}],D_{\textnormal{test}}^{i})\}_{i\in{\mathcal{I}}_{j}}$ remains the same. As a result, $\widetilde{R}$ is also fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$. In addition, the set of scores $\{v_{i}\colon i\in\widetilde{R}\}$ is fully determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, which remains the same no matter which value in $[z_{1},\dots,z_{i-1},z_{n+j},z_{i+1},\dots,z_{n}]$ the test point $Z_{n+j}$ takes on.
Recall that ^$\widehat{\mathcal{R}}_{n+j}(Y_{n+j})=\{i\in[n]:j\in\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j})\in\mathfrak{S}\}$. We let ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})$ if $j\notin\widehat{\mathcal{S}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$ and ^$\widehat{\mathcal{R}}^{+}_{n+j}=\widehat{\mathcal{R}}_{n+j}(Y_{n+j})\cup\{n+j\}$ otherwise.
We claim that given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$,
@@
where repetition of elements is allowed on both sides. In words, (34) means that no matter which value in $[z_{1},\dots,z_{n},z_{n+j}]$ is the test point $Z_{n+j}$, the left-handed side always equals $\{v_{i}\colon i\in\widetilde{R}\}$, which is determined by $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$. To see this, let $\pi$ denote the permutation on ${\mathcal{I}}_{j}$ such that $V_{i}=v_{\pi(i)}$ and $V_{n+j}=v_{\pi(n+j)}$. On this event,
@@
for all $i\in\{1,\dots,n\}$. Denoting $\widehat{\mathcal{S}}=\widehat{\mathcal{S}}^{{\textnormal{swap}(n+j,j)}}(Y_{n+j})$ for notational simplicity, we have,
$\{V_{i}\colon i\in\widehat{\mathcal{R}}_{n+j}^{+}\}$ =$\overset{\rm(a)}{=}\big{\{}V_{i}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{{\textnormal{swap}(i,j)}}(Y_{n+j})\in\mathfrak{S}\big{\}}$
=$\overset{\rm(b)}{=}\big{\{}v_{\pi(i)}\colon i\in{\mathcal{I}}_{j},j\in\widehat{\mathcal{S}}^{{\textnormal{swap}(i,j)}}(Y_{n+j}),\,\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(Y_{n+j})\in\mathfrak{S}\big{\}}$
@@
=$=\big{\{}v_{\ell}\colon\ell\in{\mathcal{I}}_{j},j\in{\mathcal{S}}\big{(}[D_{{\textnormal{calib}}}^{\ell}],D_{{\textnormal{test}}}^{\ell}\big{)},\,{\mathcal{S}}\big{(}[D_{\textnormal{calib}}^{\ell}],D_{\textnormal{test}}^{\ell}\big{)}\in\mathfrak{S}\big{\}}$
@@
Above, step (a) follows from the definition of ^$\widehat{\mathcal{R}}_{n+j}^{+}$; steps (b) and (c) are due to (LABEL:eq:representation), and step (d) follows from the definition of ${\widetilde{R}}$. We thus completed the proof of Equation (34).
Returning to the proof of (33), we have that
@@
where the last step is due to Bayes’ rule. Again, given $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, the only randomness is the assignment of $[z_{1},\dots,z_{n},z_{n+j}]$ to $(Z_{1},\dots,Z_{n},Z_{n+j})$. By (34),
@@
Since ${\widetilde{R}}$ is deterministic conditional on $[d_{j}]$ and ${\mathcal{D}}_{j}^{c}$, we know that
@@
On the other hand, it is straightforward that
Combining the above and (42), we have thus verified (LABEL:eq:equiv_cond_cov3). Then by the tower property, we conclude (33), and thus (11).
When the elements in ${\widetilde{R}}$ are distinct, we additionally have the following deterministic inequality following from the definition of the quantile function:
@@
Proof of (b). To conveniently denote repeated values in $[z_{1},\dots,z_{n},z_{n+j}]$ and $[v_{1},\dots,v_{n},v_{n+j}]$ we let $[1],[2],\dots,[n],[n+j]$ be a permutation of ${\mathcal{I}}_{j}$ so that $v_{[1]}\leq v_{[2]}\leq\cdots\leq v_{[n]}\leq v_{[n+j]}$. Then $[d_{j}]$ can be equivalently represented by $\{z_{[i]}\}_{i\in{\mathcal{I}}_{j}}$, and for any $i\in{\mathcal{I}}_{j}$,
@@
Following the proof of part (a), we know that
@@
On the other hand, due to (34) we have for any $t\in(0,1)$,
$\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}<V_{i}\}+U_{j}(1+\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}=V_{i}\})}{1+|\widehat{{\mathcal{R}}}_{n+j}(Y_{n+j})|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
$=$ $\leavevmode\nobreak\ \mathbb{P}\bigg{(}\frac{\sum_{i\in\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})}\mathds{1}\{V_{n+j}<V_{i}\}+U_{j}(\sum_{i\in\widehat{{\mathcal{R}}}_{n+j}^{+}(Y_{n+j})}\mathds{1}\{V_{n+j}=V_{i}\})}{|\widehat{{\mathcal{R}}}^{+}_{n+j}(Y_{n+j})|}\leq t,\leavevmode\nobreak\ j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{{\mathcal{S}}}\in\mathfrak{S}\,\Big{|}\,[d_{j}],{\mathcal{D}}_{j}^{c}\bigg{)}$
@@
where we used the take-out property of conditional probability/expectation in the first step; the probability in the second line is only with respect to $U_{j}\sim\textnormal{Unif}[0,1]$, and the last equality is a standard result that is straightforward to check with exchangeability (see, e.g., Vovk et al., (2005, Proposition 2.4)). Putting them together and letting $t=1-\alpha$, we have